OpenStudy (anonymous):

Solve the equation: 5x(x + 3) = 0. Step by step explanation please! Will give medal!

OpenStudy (anonymous):

I got up to 5x^2= -15x

OpenStudy (anonymous):

Ok, first you take off the parenthesis, you can do this because there is an understood one in front of them so you are just multiplying every thin by 1. Now you have 5xx + 3 = 0. You now have five x and another x so you have 6 x. Now it is 6x + 3. You subtract the three form both sides giving us 6x = -3. Just divide by six. x= - 0.5

OpenStudy (anonymous):

Im sorry i misunderstood the equation that is incorrect it should be -3

OpenStudy (anonymous):

ok.. I sorta get it...

OpenStudy (anonymous):

wait how would it become -3!

OpenStudy (anonymous):

*? not !

OpenStudy (anonymous):

5x(x + 3) = 0 5x * (x + 3) = 0 in order for it to equal zero either 5x = 0 or (x + 3) = 0 solving these gives two solutions, x = 0 or x = -3

OpenStudy (mathstudent55):

When a product equals zero, it is only possible if one of its factors equals zero. Set each factor equal to zero and solve for x. 5x(x + 3) = 0 5x = 0 or x + 3 = 0 x = 0 or x = -3